The expression (Sqrt of x + Sqrt of y)^2 is not equal to x+y for every nonnegative x and y.
1. Why is this?
2. How do I find a value of x and y for which this is true, that is (sqrt of x + sqrt of y)^2 = x + y. (Also, what is this value?)
Thank you for your time and help with this question, I hope it is clear enough to understand.)
This question is not from a text book, it is from a worksheet.

Answer by edjones(8007)

1. (sqrt(x)+sqrt(y))^2=x+2sqrt(x)sqrt(y)+y so it only equals x+y when 2sqrt(x)sqrt(y)=0
2. it will only equal 0 when x and y both equal 0.
.
Ed